# Variable Selection in Convex Piecewise Linear Regression

### v1

@@ Abstract
This paper presents Sparse Gradient Descent as a solution for variable selection in convex piecewise linear regression where the model is given as the maximum of $k$-affine functions $\bm{x}\mapsto\max_{j\in[k]}\langle\bm{a}_{j}^{\star},\bm{x}\rangle+b_{j}^{\star}$ for $j=1,\dots,k$. Here, $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ and $\{b_{j}^{\star}\}_{j=1}^{k}$ denote the ground-truth weight vectors and intercepts. A non-asymptotic local convergence analysis is provided for Sp-GD under sub-Gaussian noise when the covariate distribution satisfies sub-Gaussianity and anti-concentration property. When the model order and parameters are fixed, Sp-GD provides an $\epsilon$-accurate estimate given $\mathcal{O}(\max(\epsilon^{-2}\sigma_{z}^{2},1)s\log(d/s))$ observations where $\sigma_{z}^{2}$ denotes the noise variance. This also implies the exact parameter recovery by Sp-GD from $\mathcal{O}(s\log(d/s))$ noise-free observations. Since optimizing the squared loss for sparse max-affine is non-convex, an initialization scheme is proposed to provide a suitable initial estimate within the basin of attraction for Sp-GD, i.e. sufficiently accurate to invoke the convergence guarantees. The initialization scheme uses sparse principal component analysis to estimate the subspace spanned by $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ then applies an $r$-covering search to estimate the model parameters. A non-asymptotic analysis is presented for this initialization scheme when the covariates and noise samples follow Gaussian distributions. When the model order and parameters are fixed, this initialization scheme provides an $\epsilon$-accurate estimate given $\mathcal{O}(\epsilon^{-2}\max(\sigma_{z}^{4},\sigma_{z}^{2},1)s^{2}\log^{4}(d))$ observations. Numerical Monte Carlo results corroborate theoretical findings for Sp-GD and the initialization scheme.
Keywords— variable selection, non-linear regression, max-affine, piecewise linear
@@ 1 Introduction
We consider a multivariate regression problem where the target variable $y\in\mathbb{R}$ depends non-linearly on covariates in $\bm{x}\in\mathbb{R}^{d}$ and noise $z\in\mathbb{R}$ as
@@
where $\bm{\theta}^{\star}\in\mathbb{R}^{k(d+1)}$ collects all ground-truth parameters $\{(\bm{a}_{j}^{\star},b_{j}^{\star})\}_{j=1}^{k}\subset\mathbb{R}^{d}\times\mathbb{R}$. The max-affine structure in (2) induces a class of convex piecewise linear functions. The set of max-affine functions provides an efficient approximation of a class of smooth convex functions [Balázs et al., 2015]. It is also considered a special instance of tropical algebra [Maragos et al., 2021]. Furthermore, we assume that only up to $s$ variables in $\bm{x}$ contribute to the evaluation of each linear model in (2), i.e.
@@
where $\mathrm{supp}(\cdot)$ denotes the index set for non-zero entries of the input vector. We refer to the estimation of the ground-truth parameters in $\bm{\theta}^{\star}$ from noisy samples $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ under the sparsity constraint in (3) as sparse max-affine regression. This is equivalent to convex piecewise linear regression with variable selection. The estimation procedure naturally implements variable selection as the support of the estimate identifies the active covariate variables.
The max-affine model in (2) proved useful in various applications such as computer vision and circuit design [Magnani and Boyd, 2009]. The target variable in these applications often has a non-linear dependence on a subset of the covariates. In high dimensional settings, i.e. the number of covariates is large, identifying the set of active covariates can significantly improve the generalization performance and the interpretability of the learned model [Wasserman and Roeder, 2009]. We illustrate these aspects over data fitting to an exponential polynomial, which extends the polynomial model so that the exponents are not necessarily integers. For example, in labor economics [Zhao et al., 2016, Eq. 15], the Gross Domestic Product (GDP), $G\in\mathbb{R}$, is written as a multivariate exponential polynomial
$G=c_{1}L^{\alpha}H^{\theta}S^{\gamma}D^{\delta}+c_{2}K+c_{3}SD/K+c_{4}$ (4)
with coefficients $\{c_{i}\}_{i=1}^{4}\subset\mathbb{R}$ and exponents $\{\alpha,\theta,\gamma,\delta\}\subset\mathbb{R}$, where $L,H,S,D,K$ respectively denote labor, human capital, innovation, investment, and capital stock. Employing a $\log$-$\log$ mapping, we can transform an exponential polynomial into a sparse max-affine function [Maragos et al., 2021, Boyd et al., 2004]. The procedure is described as follows: Let
$y=\varphi\,\mathrm{log}(G),\quad\bm{x}=\varphi\begin{bmatrix}\log L,\log H,\log S,\log D,\log K\end{bmatrix}^{\mathsf{T}}.$
for some $\varphi>0$. As $\varphi$ decreases toward 00, the expression of $y$ converges to a sparse max-affine function given by
$\underset{\varphi\rightarrow 0^{+}}{\mathrm{lim}}y=\underset{j\in\{1,2,3,4\}}{\mathrm{max}}\langle\bm{\theta}_{j},[\bm{x},1]\rangle,$
where
$[\bm{\theta}_{1},\ldots,\bm{\theta}_{4}]\triangleq\begin{bmatrix}\alpha&0&0&0\\ \theta&0&0&0\\ \gamma&0&1&0\\ \delta&0&1&0\\ 0&1&-1&0\\ 0&0&0&0\end{bmatrix}.$
More generally, even when the exact dependence like (4) is not known, sparse max-affine regression can identify active variables in each component of the target-covariate dependence modeled as an exponential polynomial.
Contributions:
@@
In this paper, we present theoretical convergence guarantees for the sparse gradient descent algorithm (Sp-GD) that implements variable selection for max-affine regression. Since learning the regression parameters casts as a nonconvex optimization problem, it is important for Sp-GD to start from a suitable initialization. We propose an initialization scheme leveraging the sparsity structure backed by performance guarantees on its error bound. The theoretical results for both Sp-GD and the initialization scheme, presented respectively in pseudo Theorems 1.1 and 1.2, show that the sample complexity is governed by the number of active variables $s$ instead of the total number of variables $d$. Monte Carlo simulations in Sections 4.1 and 4.2 corroborate these theoretical guarantees.
Optimization algorithm:
@@
Sp-GD is a variant of the projected gradient descent method. Sp-GD differs from standard projected gradient descent by utilizing a generalized gradient. This gradient is an extension to non-smooth functions, such as the piecewise linear functions we are working with [Hiriart-Urruty, 1979, Clarke, 1990]. Sp-GD moves along the generalized gradient with a set of adaptive step sizes and then projects to the feasible set of sparse parameters defined in (3). Since the nonlinear least squares problem for the model in (1) is non-convex, Sp-GD provides a local convergence guarantee presented in the following pseudo-theorem.
@@
This result can be compared to a line of research on plain max-affine regression without the sparsity constraint. Ghosh et al. [2021] presented non-asymptotic convergence analyses of the alternating minimization algorithm by Magnani and Boyd [2009] under random covariates and independent stochastic noise assumptions. Later, Kim and Lee [2024] showed that stochastic gradient descent provides comparable sample complexities and estimation errors with faster convergence.
Authors Algorithm Sample complexity
Ghosh et al. [2021] Alternating Minimization $\widetilde{\mathcal{O}}(d)$
Kim and Lee [2024] First Order Methods $\widetilde{\mathcal{O}}(d)$
@@
This paper Sparse Gradient Descent $\widetilde{\mathcal{O}}\left(s\log({d}/{s})\right)$
@@
Table 1: Comparison of estimation algorithms in the sample complexity for exact parameter recovery of max-affine models from noiseless observations.
@@
Table 1 compares the required sample complexity for Sp-GD and previous parameter estimation algorithms for max-affine models. Sp-GD drops the sample complexity from $\mathcal{O}(d)$ required by previous algorithms to $\mathcal{O}(s\log(d/s))$ which is sub-linear. This marks a significant improvement especially when the number of active variables $s$ is significantly smaller than the total number of variables $d$. It is also worth noting that the convergence result in Theorem 1.1 applies to a practical implementation of SpGD. Specifically, the step sizes are given in an explicit form determined by the parameter estimates of the previous iteration. The step size for the $j$th block will concentrate around the inverse of the probability where the $j$th linear model in (2) achieves the maximum. In contrast, Kim and Lee [2024] used an unspecified constant step size to prove the local convergence of their first-order methods. Our step size strategy also applies to plain max-affine regression.
Initialization:
@@
Recall that Theorem 1.1 requires a suitable initial estimate to guarantee local convergence of Sp-GD. To obtain this desired initialization, one may use the spectral method for max-affine regression [Ghosh et al., 2021]. Simply stated, their initialization scheme first employs principal component analysis (PCA) to estimate the span of $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$. Then a discrete search over an $r$-covering in the span of the principal components is applied. It requires $\mathcal{O}(\epsilon^{-2}d\log^{3}d)$ observations for their method to guarantee an $\epsilon$-accurate estimate with $r=\mathcal{O}(\epsilon)$ for fixed $\bm{\theta}^{\star}$ and noise level. This voids the gain due to Theorem 1.1. We propose a variant of the initialization by Ghosh et al. [2021] that substitutes PCA with sparse PCA (sPCA) and provides the desired initial estimate from fewer observations in the following special scenario: Suppose that the sparse coefficient vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ are simultaneously supported within a set of cardinality $s$, i.e.
@@
Compared to previous work [Ghosh et al., 2021], the initialization scheme leveraging the joint sparsity drops the linear dependence on the ambient dimension $d$ and replaces it with a quadratic dependence on $s$. Our implementation and theoretical analysis build on sPCA by semi-definite program by Vu et al. [2013] where the estimation error decays as $\mathcal{O}(\sqrt{s^{2}\log d/n})$. A recent framework improved the error bound to $\mathcal{O}(\sqrt{s\log d/n})$ under the assumption that data are obtained through a linear transform acting on special multivariate distributions [Wang et al., 2014]. However, this assumption is not satisfied by data generated with the max-affine model.
Related work:
Variable selection for linear regression models has been extensively studied in the literature (e.g. [George, 2000, Wasserman and Roeder, 2009]). Various optimization approaches including Lasso and basis pursuit have been studied [Hastie et al., 2009, Wainwright, 2019]. In practical applications, the covariate-target dependence often exhibits non-linearity not captured by linear models. There have been efforts to extend variable selection for linear regression to the non-linear models via basis change [Meier et al., 2009, Koltchinskii and Yuan, 2010]. They modeled the nonlinear function describing the target-covariate relation as a sparse superposition over a certain function basis (e.g. splines). Variable selection in their methods identifies active multivariate functions in the basis expansion. This approach is similar to the well-known order selection for polynomial regression [Peixoto, 1987]. Unlike the above model, similar to variable selection for linear regression, the sparse max-affine model directly identifies the active covariate variables to explain the target variable.
Variable selection has also been used along with dimensionality reduction for high-dimensional nonlinear models. The sparse minimum average variance estimator (sMAVE) [Wang and Yin, 2008] and sparse sliced inverse regression (sSIR) [Tan et al., 2020] considered the model where the target is given by a nonlinear function acting on linear combinations of a subset of covariates. Both sMAVE and sSIR estimate the span of the weight vectors in the linear combinations without attempting to learn the unknown non-linear function. Only asymptotic analyses of these algorithms exist in the high sample regime, i.e. the total number of samples goes to infinity while the model order remains fixed. The sparse max-affine model corresponds to a special case of their model where the non-linear function is known. Under this scenario, we provide a stronger non-asymptotic theoretical analysis for variable selection for the nonlinear max-affine model.
Paper organization:
@@
Section 2 presents Sp-GD and its non-asymptotic theoretical guarantees as local analysis of sparse max-affine regression. Section 3.1 describes the initialization scheme the corresponding non-asymptotic theoretical guarantees. Section 4 presents the numerical results that corroborate the theoretical guarantees for Sp-GD and the initialization scheme. Section 5 provides final remarks and future directions.
Notation:
[…]

### v2

# Sparse Max-Affine Regression

@@ Abstract
This paper presents Sparse Gradient Descent as a solution for variable selection in convex piecewise linear regression where the model is given as the maximum of $k$-affine functions $\bm{x}\mapsto\max_{j\in[k]}\langle\bm{a}_{j}^{\star},\bm{x}\rangle+b_{j}^{\star}$ for $j=1,\dots,k$. Here, $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ and $\{b_{j}^{\star}\}_{j=1}^{k}$ denote the ground-truth weight vectors and intercepts. A non-asymptotic local convergence analysis is provided for Sp-GD under sub-Gaussian noise when the covariate distribution satisfies the sub-Gaussianity and anti-concentration properties. When the model order and parameters are fixed, Sp-GD provides an $\epsilon$-accurate estimate given $\mathcal{O}(\max(\epsilon^{-2}\sigma_{z}^{2},1)s\log(d/s))$ observations where $\sigma_{z}^{2}$ denotes the noise variance. This also implies the exact parameter recovery by Sp-GD from $\mathcal{O}(s\log(d/s))$ noise-free observations. Since optimizing the squared loss for sparse max-affine is non-convex, an initialization scheme is proposed to provide a suitable initial estimate within the basin of attraction for Sp-GD, i.e. sufficiently accurate to invoke the convergence guarantees. The initialization scheme uses sparse principal component analysis to estimate the subspace spanned by $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$, then applies an $r$-covering search to estimate the model parameters. A non-asymptotic analysis is presented for this initialization scheme when the covariates and noise samples follow Gaussian distributions. When the model order and parameters are fixed, this initialization scheme provides an $\epsilon$-accurate estimate given $\mathcal{O}(\epsilon^{-2}\max(\sigma_{z}^{4},\sigma_{z}^{2},1)s^{2}\log^{4}(d))$ observations. A new transformation named Real Maslov Dequantization (RMD) is proposed to transform sparse generalized polynomials into sparse max-affine models. The error decay rate of RMD is shown to be exponentially small in its temperature parameter. Furthermore, theoretical guarantees for Sp-GD are extended to the bounded noise model induced by RMD. Numerical Monte Carlo results corroborate theoretical findings for Sp-GD and the initialization scheme.
Keywords— V ariable selection, nonlinear regression, convex regression, piecewise linear, generalized polynomial, posynomial.
@@ 1 Introduction
We consider a multivariate regression problem where the target variable $y\in\mathbb{R}$ depends nonlinearly on covariates in $\bm{x}\in\mathbb{R}^{d}$, and noise $z\in\mathbb{R}$ as
@@
where $\bm{\theta}^{\star}\in\mathbb{R}^{k(d+1)}$ collects all ground-truth parameters $\{(\bm{a}_{j}^{\star},b_{j}^{\star})\}_{j=1}^{k}\subset\mathbb{R}^{d}\times\mathbb{R}$. The max-affine structure in (2) induces a class of convex piecewise linear functions. The set of max-affine functions provides an efficient approximation of a class of smooth convex functions (Balázs et al., 2015). It is also considered a special instance of tropical algebra (Maragos et al., 2021). Furthermore, we assume that only up to $s$ variables in $\bm{x}$ contribute to the evaluation of each linear model in (2), i.e.
@@
where $\mathrm{supp}(\cdot)$ denotes the index set for non-zero entries of the input vector. We denote such models as sparse max-affine. We refer to the estimation of the ground-truth parameters in $\bm{\theta}^{\star}$ from noisy samples $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ under the sparsity constraint in (3) as sparse max-affine regression. This is equivalent to convex piecewise linear regression with variable selection. The estimation procedure naturally implements variable selection as the support of the estimate identifies the active covariate variables.
Motivation
Estimating nonlinear functions under convexity constraints, known as convex regression, constitutes a fundamental problem across multiple disciplines, including econometrics (Merton and Samuelson, 1992), geometric programming (Magnani and Boyd, 2009), and reinforcement learning (Hannah et al., 2014). Additional applications of convex regression can be found in circuit design (Hannah and Dunson, 2012) and queuing theory (Chen and Yao, 2001). Known approaches for convex regression can be categorized by their key assumptions on the true underlying function $f:\mathbb{R}^{d}\to\mathbb{R}$ as follows:
  1. 1.
    Non-parametric: $f$ satisfies minimal conditions like continuity or smoothness.
  2. 2.
    Semi-parametric: $f$ is expressed as $\bm{x}\mapsto f(\bm{x})=g(\bm{U}\bm{x})$, with unknown $\bm{U}\in\mathbb{R}^{K\times d}$ and unknown nonlinear map $g:\mathbb{R}^{K}\to\mathbb{R}$. This is referred to as the multi-index model if $K>1$ or the single-index model if $K=1$.
  3. 3.
    Parametric: $f$ has a known form involving a few unknown parameters.
Each approach involves trade-offs between its advantages and limitations. Training methods for non-parametric regression typically solve a quadratic program with $\mathcal{O}(n)$ inequality constraints, where $n$ denotes the number of samples (Balázs et al., 2015; Hannah and Dunson, 2013). While non-parametric convex regression offers broad applicability, it suffers from exponentially increasing sample complexity on the order of $\mathcal{O}(e^{d})$ in the number of covariates $d$ and incurs a high computational cost of $\mathcal{O}(\mathrm{poly}(nd))$. In contrast, a semi-parametric method provided a sharper result on a further restricted convex single-index model with the generalization error rate of $\mathcal{O}\left(\frac{d^{2/5}}{n^{2/5}}\right)$ (Kuchibhotla et al., 2023). With the same goal, parametric methods employed an alternative compact model composed of piecewise linear functions with a fixed number of linear components, which can be considered as a special case of multi-index models with $g$ fixed to the max function. Computationally efficient solutions such as alternating partitioning and minimization (Magnani and Boyd, 2009) and first-order methods (Kim and Lee, 2024) have been proposed. When the number of linear components is fixed, and under certain covariate conditions that generalize the standard Gaussian model, the parametric method can achieve a computational cost of $\mathcal{O}(nd)$, with the required sample size $n$ scaling linearly as $\mathcal{O}(d)$ (Ghosh et al., 2021; Kim and Lee, 2024).
In many real-world problems, the outcome of interest is often driven by a small subset of key input factors, whose effects can be highly nonlinear. Variable selection seeks to identify this unknown subset of active inputs by promoting sparsity in the regression coefficients. In the high-dimensional settings, where the number of covariates $d$ is large, variable selection enhances the generalization performance and improves the interpretability of the learned model (Wasserman and Roeder, 2009). However, establishing theoretical performance guarantees for variable selection in general nonlinear regression models is widely recognized in the literature as a significant challenge (Lafferty and Wasserman, 2008; Bertin and Lecué, 2008). In particular, existing theoretical results are limited and focus on special restrictive cases. For example, under the smoothness of the nonlinear function given by a norm induced by the Fourier series, asymptotic consistency for support recovery can be achieved with $n=\mathcal{O}(e^{s}\log d)$ by a non-parametric method (Comminges and Dalalyan, 2012), where $s$ is the number of relevant variables. Regarding the semi-parametric approach, a generalization error rate decaying as $\mathcal{O}(\frac{s\log d}{n})$ has been established for the single-index model (Alquier and Biau, 2013; Radchenko, 2015). However, for the multi-index model ($K>1$), existing semi-parametric results are limited to establishing asymptotic consistency for support recovery (Wang et al., 2015). For the parametric approach, the parameter estimation error rate of $O\left(\frac{s\log d}{n}\right)$ is attainable under the stringent assumption of a known monotone single-index model (Yang et al., 2016). Variable selection has been studied for deep neural networks, which are also parametric models, and support recovery is shown to be asymptotically consistent (Yang et al., 2024). In summary, variable selection improves the dependence in sample complexity on $d$ by substituting it with $s$, where $s\ll d$, up to an extra logarithmic factor in $d$.
Convex regression is another instance of nonlinear regression where variable selection has been shown to provably improve the generalization performance. A seminal work by Xu et al. (Xu et al., 2016) showed that variable selection in convex non-parametric regression can be achieved with $n=\mathcal{O}(s^{7}\log^{3}d)$ under the assumption that the true underlying function is expressed as a superposition of univariate convex functions. This paper presents a variable selection scheme for convex regression on a more flexible subclass of convex functions at a lower sample complexity of $n=\mathcal{O}(s\log d)$. The max-affine function in (2) indeed provides a parametric model for convex piecewise linear multivariate functions. As we will discuss later, the max-affine model is known to be equivalent to a convex ReLU neural network (NN) (Zhang et al., 2018). The approximation power of ReLU-NN is well studied in the literature (Schmidt-hieber, 2020). To the author’s best knowledge, this is the first work to consider a parametric approach to sparse convex regression. The sample complexities of convex regression methodologies with and without variable selection are summarized in Table 1.
Variable selection Assumptions other than convexity Sample complexity
Balázs et al. (2015) continuous & bounded gradients $\mathcal{O}(e^{d})$
Xu et al. (2016) superposition of univariate functions $\mathcal{O}(e^{s}\log d)$ [^1] Xu et al. also show that the support recovery can be done with $n=\mathcal{O}(s^{7}\log^{3}d)$.
Kuchibhotla et al. (2023) single-index model $\mathcal{O}(d)$
Ghosh et al. (2021) piecewise linear $\mathcal{O}(d)$
This paper piecewise linear $\bm{\mathcal{O}(s\log d)}$
Table 1: Comparison of methodologies for convex regression with associated assumptions on the true convex function and non-asymptotic sample complexities.
Next, we delve into further detail regarding the connection between sparse max-affine models and deep ReLU-NNs. The work by (Zhang et al., 2018) shows that a max-affine model is equivalent to an $L$-layer ReLU-NN with nonnegative hidden layer weights. Furthermore, (Zhang et al., 2018, Theorem 5.4) states that any max-affine model with order $k$ can be written as an $L$-layer ReLU-NN with $L\leq 2+\lceil\log_{2}k\rceil$. Therefore, training a sparse max-affine model is equivalent in spirit to ReLU-NN training with a regularization term that enforces sparsity (e.g., an $\ell_{1}$ penalty term on the network weights). Variable selection techniques for NN are employed to handle high-dimensional datasets with a small number of samples (see (Yang et al., 2024) for a recent review). However, non-asymptotic theoretical guarantees for NN are limited to restrictive cases such as the two-layer model in (Li et al., 2020) without sparsity. Variable selection in NN training is shown to be consistent only when the number of samples is infinitely large (Yang et al., 2024, Theorem 4.2). Furthermore, the sparse max-affine model provides better interpretability as the contributing weights of each covariate are clearly shown. We also note that the sparse max-affine model is a natural object from tropical algebra called the tropical rational functional. Therefore, compared to ReLU-NN, sparse max-affine models inherit the existing geometric results from tropical algebra (Zhang et al., 2018).
Another main motivation for studying the sparse max-affine model is that it provides an efficient approximation to a broad class of generalized sparse polynomials (Jameson, 2006). This model is written as
$w=g(u_{1},\dots,u_{d})=\sum_{j=1}^{k}c_{j}^{\star}\left(\prod_{\ell\in\mathcal{I}_{j}}u_{l}^{\alpha^{\star}_{j,\ell}}\right),$ (4)
with real-valued exponents $\{\alpha^{\star}_{j,\ell}\}_{\ell\in\mathcal{I}_{j}}\subset\mathbb{R}$ instead of the integer-valued exponents in a sparse polynomial, a real-valued coefficient $c_{j}$, and the active monomial indices $\mathcal{I}_{j}\subseteq[d]$ for all $j\in[k]$. Variable selection here refers to finding $\{\mathcal{I}_{j}\}_{j=1}^{k}$. When the coefficients $c_{j}$’s are positive, (4) reduces to a posynomial (a special case of the generalized polynomial). Employing a $\log$-$\log$ mapping, called Maslov dequantization, one can transform a sparse posynomial into a sparse max-affine function (Maragos et al., 2021; Boyd et al., 2004). Even with the positivity constraint on the weights $c_{j}$’s, the model by (4) is not restricted to a convex function, but does not provide sufficient flexibility to model complex relations in real-world applications. Furthermore, there has been no non-asymptotic approximation analysis for the Maslov dequantization, even for sparse posynomials. This paper will also further extend this approach so that a sparse max-affine model can approximate the generalized sparse polynomial model in (4) without requiring the positivity constraint backed by a non-asymptotic error bound.
Contributions
@@
In this paper, we present theoretical convergence guarantees for the sparse gradient descent algorithm (Sp-GD) that implements variable selection for max-affine regression. Since learning the regression parameters is cast as a nonconvex optimization problem, it is important for Sp-GD to start from a suitable initialization. We propose an initialization scheme leveraging the sparsity structure backed by performance guarantees on its error bound. The theoretical results for both Sp-GD and the initialization scheme, presented respectively in pseudo Theorems 1.1 and 1.2, show that the sample complexity is governed by the number of active variables $s$ instead of the total number of variables $d$. Furthermore, we propose the Real Maslov dequantization to approximate any generalized sparse polynomial via a sparse max-affine function. The quantization error of this approximation is presented in pseudo Theorem 1.3. We also extend the theoretical convergence guarantees of Sp-GD to this context in Section 4. Monte Carlo simulations in Sections 5.1 and 5.2 corroborate these theoretical guarantees.
Optimization algorithm
@@
Sp-GD is a variant of the projected gradient descent method. Sp-GD differs from standard projected gradient descent by utilizing a generalized gradient. This gradient is an extension to non-smooth functions, such as the piecewise linear functions we are working with (Hiriart-Urruty, 1979; Clarke, 1990). Sp-GD moves along the generalized gradient with a set of adaptive step sizes and then projects to the feasible set of sparse parameters defined in (3). Since the nonlinear least squares problem for the model in (1) is non-convex, Sp-GD provides a local convergence guarantee presented in the following pseudo-theorem.
@@
This result can be compared to a line of research on plain max-affine regression without the sparsity constraint. The authors in (Ghosh et al., 2021) presented non-asymptotic convergence analyses of the alternating minimization algorithm by (Magnani and Boyd, 2009) under random covariates and independent stochastic noise assumptions. Later, the authors in (Kim and Lee, 2024) showed that stochastic gradient descent provides comparable sample complexities and estimation errors with faster convergence.
Algorithm Sample complexity Step-size for theoretical guarantees
Ghosh et al. (2021) alternating minimization ${\mathcal{O}}(d)$ NA
Kim and Lee (2024) first order methods ${\mathcal{O}}(d)$ unspecified constant
@@
This paper sparse gradient descent $\bm{\mathcal{O}\left(s\log d\right)}$ adaptive formula
@@
Table 2: Comparison of max-affine regression algorithms in the sample complexity for exact parameter recovery from noiseless observations.
@@
Table 2 compares the required sample complexity for Sp-GD and previous parameter estimation algorithms for max-affine models. Sp-GD drops the sample complexity from $\widetilde{\mathcal{O}}(d)$ required by previous algorithms to $\widetilde{\mathcal{O}}(s\log(d/s))$ which is sub-linear. This marks a significant improvement, especially when the number of active variables $s$ is significantly smaller than the total number of variables $d$. It is also worth noting that the convergence result in Theorem 1.1 applies to a practical implementation of Sp-GD. Specifically, the step sizes are given in an explicit form determined by the parameter estimates of the previous iteration. The step size for the $j$th block will concentrate around the inverse of the probability where the $j$th linear model in (2) achieves the maximum. In contrast, the authors in (Kim and Lee, 2024) used an unspecified constant step size to prove the local convergence of their first-order methods. Therefore, our step size strategy improves the max-affine regression strategy in (Kim and Lee, 2024) even in the non-sparse case.
Initialization
@@
Recall that Theorem 1.1 requires a suitable initial estimate to guarantee local convergence of Sp-GD. To obtain this desired initialization, one may use the spectral method for max-affine regression presented by the authors in (Ghosh et al., 2021). Simply stated, their initialization scheme first employs principal component analysis (PCA) to estimate the span of $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$. Then a discrete search over an $r$-covering in the span of the principal components is applied. It requires $\mathcal{O}(\epsilon^{-2}d\log^{3}d)$ observations for their method to guarantee an $\epsilon$-accurate estimate with $r=\mathcal{O}(\epsilon)$ for fixed $\bm{\theta}^{\star}$ and noise level. This voids the gain due to Theorem 1.1. We propose a variant of the initialization by (Ghosh et al., 2021) that substitutes PCA with sparse PCA (sPCA) and provides the desired initial estimate from fewer observations in the following special scenario: Suppose that the sparse coefficient vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ are simultaneously supported within a set of cardinality $s$, i.e.
@@
Compared to previous work (Ghosh et al., 2021), the initialization scheme leveraging the joint sparsity drops the linear dependence on the ambient dimension $d$ and replaces it with a quadratic dependence on $s$. Our implementation and theoretical analysis build on sPCA as a semi-definite program by (Vu et al., 2013) where the estimation error decays as $\mathcal{O}(\sqrt{s^{2}\log d/n})$. A recent framework improved the error bound to $\mathcal{O}(\sqrt{s\log d/n})$ under the assumption that data are obtained through a linear transform acting on special multivariate distributions (Wang et al., 2014). However, this assumption is not satisfied by data generated with the max-affine model.
Approximating generalized polynomials
We overcome the main limitation of Maslov dequantization by extending this transformation to real coefficients in (4) via the Real Maslov dequantization (RMD), which is written as
$y=\mathrm{Re}\{\varsigma\log w\},\quad x_{l}=\varsigma\log u_{l},\quad\forall l\in[d]$ (6)
where $\varsigma>0$ is a chosen temperature parameter and $\log(re^{i\theta})=\log r+i\mathrm{mod}(\theta,2\pi)$. We will use these definitions to state the following theorem on the approximation error of RMD.
Theorem 1.3 (Informal).
Consider the generalized polynomial model in (4) and its transformation via RMD in (6) for some $\varsigma>0$. Collect the transformed covariates in $\bm{x}=[x_{1};\ldots;x_{d}]$. The transformed target can be written as
$y=\max_{j\in[k]}\langle\bm{\theta}_{j}^{\star},[\bm{x};1]\rangle+z_{\varsigma},$
where $\bm{\theta}_{j}^{\star}=[\alpha_{j,1}^{\star};\ldots;\alpha_{j,d}^{\star};\varsigma\log|c_{j}^{\star}|]$ and $z_{\varsigma}$ is the dequantization error which decays exponentially in $\varsigma$, i.e. $z_{\varsigma}\leq c\exp(\varsigma^{-1})$. Consequently, when $\varsigma\to 0$, RMD reduces to a sparse-max affine model.
The class of models in (4) is very rich and has been applied to various contexts in finance and economics. For example, in labor economics (Zhao et al., 2016, Eq. 15), the Gross Domestic Product (GDP), $G\in\mathbb{R}$, is written as a multivariate generalized polynomial
$G=c_{1}L^{\alpha}H^{\theta}S^{\gamma}D^{\delta}+c_{2}K+c_{3}SD/K+c_{4}$ (7)
with coefficients $\{c_{i}\}_{i=1}^{4}\subset\mathbb{R}$ and exponents $\{\alpha,\theta,\gamma,\delta\}\subset\mathbb{R}$, where $L,H,S,D,K$ respectively denote labor, human capital, innovation, investment, and capital stock. Using RMD as defined in (6) with $\varsigma\to 0^{+}$, we get
$y=\underset{j\in[4]}{\mathrm{max}}\langle\bm{\theta}_{j},[\bm{x},1]\rangle,$
where
$\bm{\theta}_{1},\ldots,\bm{\theta}_{4}]\triangleq\begin{bmatrix}\alpha&0&0&0\\ \theta&0&0&0\\ \gamma&0&1&0\\ \delta&0&1&0\\ 0&1&-1&0\\ 0&0&0&0\end{bmatrix}$,
which is clearly a sparse max-affine model. Another example of such modeling is the McCallum Gravity Equation (Anderson and Van Wincoop, 2003) of party ”11”, written as
$E_{1}=\sum_{i\neq 1}c_{i}G_{1}^{\alpha}G_{i}^{\beta}D_{1\to i}^{\delta},$
where $E_{1}$ denotes the exports of party $1$, $G_{i}$ is the GDP of the $i$th party, and $D_{1\to i}$ is the distance between parties $1$ and $i$. Furthermore, we can also find applications of the generalized polynomial models in other fields, such as differential equation modeling for fluid mechanics (Ranganathan and Minchew, 2024), characteristic equations for permeability (Siddiqui et al., 2008), and power control in cellular systems (Chiang et al., 2017). Finally, consider the generalized rational function that is the division of two models from (4). Applying RMD to this rational function leads to a difference of max-affine model which we are currently investigating in an independent work.
Paper organization
@@
Section 2 presents Sp-GD and its non-asymptotic theoretical guarantees as a local analysis of sparse max-affine regression. Section 3.1 describes the initialization scheme and the corresponding non-asymptotic theoretical guarantees. Section 4 presents the theoretical guarantees for the dequantization error by RMD for generalized sparse polynomials and the convergence guarantee for Sp-GD in this context. Section 5 presents the numerical results that corroborate the theoretical guarantees for Sp-GD and the initialization scheme. Section 6 provides final remarks and future directions.
Notation
We use lightface characters to denote scalars, lowercase boldface characters to denote column vectors and uppercase boldface to denote matrices. We also adopt the symbols for the max and min operators in the lattice theory, i.e. $a\vee b=\max(a,b)$ and $a\wedge b=\min(a,b)$ for $a,b\in\mathbb{R}$. We use multiple matrix norms. The Frobenius norm, the spectral norm, and the largest magnitude of entries will be respectively denoted by ${\left\lVert\cdot\right\rVert}_{\mathrm{F}}$, $\left\lVert\cdot\right\rVert$}, and $\left\lVert\cdot\right\rVert$$_{\infty}$. We use a shorthand notation $[d]$ for the set $\{1,2,\dots,d\}$. For a column vector $\bm{x}\in\mathbb{R}^{d}$, its sub-vector with the entries indexed by $\mathcal{S}\subset[d]$ is denoted by $\bm{x}]_{\mathcal{S}}$. Similarly, for a matrix $\bm{X}\in\mathbb{R}^{d\times d}$, its submatrix with the entries indexed by $\mathcal{S}_{1}\times\mathcal{S}_{2}\subset[d]\times[d]$ is denoted by $\bm{X}]_{\mathcal{S}_{1},\mathcal{S}_{2}}$. Finally, we denote by $C,C_{1},C_{2},\dots$ universally absolute constants, not necessarily the same at each occurrence.